An object has a constant acceleration a=4m/s². It's velocity is 1m/s at t=0, when it is at x=7m. How fast is it moving when it is at x=8m? At what time is this?
1
Expert's answer
2020-03-05T09:34:15-0500

"a=4\ m/s^2"

"v_i=1\ m/s, \ \ x_i=7\ m" , "t=0 \ s"

"v_f \ - ?, \ x_f=8\ m" , "t - ?"


Solution:

Using formula "v_f^2=v_i^2 +2ad, (d" - displacement")" we have:

"v_f^2=1^2+2\times 4\times (8-7)=9, \; \; v_f=3\ m/s"

To find time (when "v_f=3\ m/s" ) we can use formula for velocity of object with constant acceleration:

"v_f=v_i+at,"

"t=\frac{v_f-v_i}{a}=\frac{3-1}{4}=0.5 \ s"

Answer: "v_f=3 \ m/s, \; t=0.5\ s."